OpenStudy (anonymous):

1 + cos(2x) = cot(x)sin(2x)

OpenStudy (raden):

use the identity : cos(2x) = 2cos^2 (x) - 1 so, the left side becomes 1 + cos(2x) = 1 + 2cos^2 (x) - 1 = 2cos^2 (x) then use the identity to simplify in right side, they are cot(x) = cos(x)/sin(x) and sin(2x) = 2sin(x)cos(x) therefore, cot(x)sin(2x) = cos(x)/sin(x) * 2sin(x)cos(x) = 2cos^2 (x) LHS = RHS QED